A pilot sets plane's speed for 500km/h. If they want the plane to head due south, but there is an 80km/h blowing from the west, then what should they set as the plane's heading?
1
Expert's answer
2019-03-22T10:39:08-0400

Plane's heading should be south-west at an angle (between the plane's course and the south)

"\alpha = \arcsin{\frac{v_w}{v_p}} = \arcsin{\frac{80}{500}} \approx 9.21^\circ."